# A Plug-and-Play Method for Guided Multi-contrast MRI Reconstruction based on Content/Style Modeling

### v4

@@ 4.1.4 Preliminary radiological evaluation
Finally, as a supplement to the LUMC benchmark, we conducted a preliminary radiological evaluation with the goal of assessing the quality of PnP-CoSMo reconstructions in a concrete clinical context. The study was conducted on a small sample of the LUMC-TRA recon-test set. First, all 20 subjects from this set (see Table 2) were screened by a junior radiologist who selected three best cases with prominently visible brain metastases. The scoring was then performed independently by two senior neuroradiologists. The scoring system comprised two sets of criteria, namely visual quality and pathology. Four visual quality criteria were used, namely sharpness, noise, artifacts, and tissue contrast between gray and white matter and CSF. For pathology, we used three criteria, namely the number and sharpness of hyperintense areas within or surrounding metastases and the overall diagnostic quality of the scan for brain metastases. The images were scored using a five-point Likert scale (Mason et al., 2019) – (1) non-diagnostic, (2) poor, (3) fair, (4) good, and (5) excellent diagnostic quality. Furthermore, we used PnP-CNN as a non-guided plug-and-play reconstruction baseline. We evaluated the two methods at four clinically realistic accelerations of , which are close to the routine clinical acceleration of =2 for this sequence. In total, nine images per patient, including the (24=8) reconstructions and the clinical ground truth (acquired at =2 and reconstructed by the vendor software), were presented to the two observers who scored each image individually, blinded to the reconstruction method and . The scores were then processed as follows. Given a reconstruction method and , the scores for each Likert-scale criterion were expressed as an aggregate by computing the mean across the three subjects and the two observers. The number of visible metastases was expressed as the fraction of total metastases across the three subjects. Moreover, the agreement between the two observers was measured using Spearman’s rank correlation coefficient, computed separately on the Likert-scale scores and the metastasis count.
@@ 4.2 Implementation details
We used a similar general residual architecture for content/style encoders and decoders as in the original MUNIT paper (Xuan et al., 2022), except with an additional layer at the end of the content encoder to produce content maps of the specified number of channels. We used 2 or 4 content channels, depending on the content downsampling factor. The discriminators were implemented as multi-scale PatchGAN networks enabling them to locally assess the input images for realism at different scales. We used 1 scale in the BrainWeb simulations and 3 scales in the NYU and LUMC benchmark. We additionally conditioned the discriminators on foreground masks of the images to penalize background signal, and we used spectral normalization (Miyato et al., 2018) to stabilize the GAN training. For hyperparameters, we found that pre-training loss weight values and paired fine-tuning values worked generally well for the in vivo datasets. In the benchmarks, the step size of the CR module of PnP-CoSMo was tuned per acceleration for each dataset and the number of iterations was set to 200.
In the simulations, all content/style models were pre-trained for 200k and were not fine-tuned for the sake of simplicity, unless stated otherwise in Section 5.1. In both NYU and LUMC benchmarks, the content/style models were pre-trained for 400k iterations and fine-tuned for 50k iterations. We empirically found with a content capacity =1 of the content/style model as optimal for NYU dataset, whereas models with capacity =1/4 were optimal for the LUMC datasets. All software was implemented in PyTorch, and image registration was performed using Elastix (Klein et al., 2009). All training runs were performed on a compute node with an NVIDIA Quadro RTX 6000 GPU.
@@ 5.1.2 Content capacity analysis
PnP-CoSMo relies on a reference image of sufficiently high resolution to provide guidance to the reconstruction. The goal of this experiment was to determine the effect of lowering the reference image resolution on the reconstruction quality and using the content/style model’s content capacity (Eq. (14)) to explain this effect. We simulated 4 datasets, denoted as RefRes-, where represents the reference domain downsampling factor. In the =1 case, the T1W/T2W images had the same resolution (as that of the underlying tissue maps), whereas in the subsequent cases, T1W images were blurred to contain only the lower frequency components, while maintaining the same spatial resolution. For each dataset, we trained 3 content/style models with content capacity =, where is the content downsampling factor, which depends on the number of up/downsampling blocks in the networks. E.g. the =(1/2)2 model produced content maps half the spatial resolution of the =1 model.
@@
5.1.3 Convergence
We explored the convergence of PnP-CoSMo by comparing it with two variants. The first used true content of the ground truth T2W image, thus assuming zero content discrepancy and representing an upperbound of PnP-CoSMo. The second used reference content with CR module disabled, representing a lowerbound of PnP-CoSMo where the non-zero content discrepancy is not corrected. Here and in the following simulation experiments, we used the RefRes-1 dataset and the =1 model which was additionally fine-tuned on 2 training volumes for 50k iterations. Fig. 4 shows the convergence curves and intermediate reconstructions. The upperbound version, given , converged in a single iteration, while with , the ablated version converged equally fast but to a sub-optimal solution. Enabling the CR module closed the gap with the upperbound, although at the cost of slower convergence. It is also observed that the content discrepancy reduced when using CR and that the style estimate always converged within the first few iterations, as we hypothesized in Section 3.3.
5.1.4 Effectiveness in resolving contrast-specific structure
In order to test the effectiveness of the CR module in resolving structure present in the target contrast but absent in the reference contrast, we simulated a lesion in the T2W image and performed reconstruction with and without the CR module. As shown in Fig. 5, the CR module was able to recover the lesion fully at =2 and substantially at =4.
5.1.5 Sensitivity to misalignment of the reference
@@
5.1.6 Influence of the CR step size
The hyperparameter in Algorithm 1 controls the CR step size. Fig. 7 shows reconstruction quality as a function of acceleration and k-space noise level, where we observe that the optimal decreased with the amount and quality of k-space data. This was expected as enforcing agreement with the k-space gradually becomes less beneficial, thereby contributing less to reconstruction quality.
@@ 5.2 Benchmark on public NYU fastMRI DICOM data
In the NYU benchmarking setup considered here, let the T1W-guided T2W reconstruction task be denoted as Task-1 and the T2W-guided T1W reconstruction task as Task-2. Fig. 8 shows evaluation metrics for PnP-CoSMo and the end-to-end baselines on these two guided reconstruction tasks. We observe three main trends. First, the end-to-end methods consistently dropped in performance on the out-of-distribution tasks (0.05 for all 3 models on both tasks at all 4 accelerations). Second, PnP-CoSMo outperformed all models on out-of-distribution tasks (0.05 throughout). Since PnP-CoSMo does not depend on any task-specific k-space training data owing to both its plug-and-play design and because of the bi-directionality of the content/style model it uses, it is not susceptible to the distribution shift of this kind. Third, considering the in-distribution model performance, PnP-CoSMo largely outperformed MoDL (0.05 throughout, except SSIM at =2 and DISTS at on Task-1, and DISTS at =2 on Task-2) and MTrans (0.05 throughout), and was roughly comparable to MC-VarNet (e.g. with 0.05 on Task-2 in terms of DISTS at =3 and SSIM at =4) while often outperforming it (0.05 throughout, except SSIM at and DISTS at on Task-1, and SSIM at and DISTS at =2 on Task-2). As shown in Figs. 9 for Task-2 and 10a for Task-1, PnP-CoSMo could resolve certain fine details even at the high acceleration of =5, which the end-to-end baselines failed at. In the pathological case shown in Fig. 10b, PnP-CoSMo produced sharper reconstructions while preserving the structure of the edema better than (the in-distribution) MC-VarNet, despite slightly lower metrics, when many features of the edema were absent in the reference image.
@@ 5.4 Preliminary radiological evaluation
Fig. 13 plots the aggregated scores of our preliminary radiological evaluation and Fig. 14 shows a set of representative example reconstructions. At , PnP-CoSMo produced reconstructions of at least ”fairly diagnostic” level (Likert score 3) of image sharpness, artifacts, tissue contrast, and overall diagnostic quality, while enabling the detection of at least 80% of all metastases. In terms of artifacts, PnP-CoSMo was consistently superior, as seen also in Fig. 14 (blue arrows). In terms of tissue contrast, it matched or exceeded PnP-CNN at and produced sharper reconstructions at (purple arrows in Fig. 14). However, in terms of noise, it was equal or slightly worse. The may be explained by the difference in nature between the denoiser-type model of PnP-CNN and the more complex content/style model of PnP-CoSMo, potentially leading to noise-like grainy texture in the latter which also interferes with image sharpness and tissue contrast. In terms of the pathology criteria, while both algorithms scored similarly on the metastasis count, PnP-CoSMo had slightly better diagnostic quality at =6. The general difference in reconstruction quality at =6 can be seen in Fig. 14 where, compared to PnP-CoSMo, the PnP-CNN reconstruction shows severe reconstruction artifacts, blurrier and more distorted features (purple arrows), and a poorly visible hyperintense area of a brain metastasis (salmon-pink arrows).
@@ 6 Discussion
Optimal content capacity (Section 3.2.3) was defined as a quantity representing the amount of shared local information contained in a two-contrast dataset. In terms of this quantity, we analyzed the effect of reference image resolution on reconstruction quality, obtaining a lowerbound for PnP-CoSMo to be effective (Section 5.1.2), specifically a reference resolution 1/8 of the target image resolution, corresponding to =1/16 of the total spatial information in the target. This quantity would depend on more fundamental factors such as MR sequence types, which we did not investigate here. While we limited our experiments to T1W and T2W sequences, PnP-CoSMo is, in principle, applicable to any pair of contrasts, with the image quality limited mainly by the amount of structural information shared between them. Future work could help empirically determine which sequence pairs are more amenable to guided reconstruction than others. A consequence of our MUNIT-based model design was the model’s lack of dependence on the uni-directionality of the guided reconstruction problem, which allowed greater generalizability by design (Section 5.2). It is, in principle, also possible to model the content and style for more than two contrasts. This would enable not only the generalizability of our approach to any combinatorial set of guided reconstruction tasks, but also better estimation of the target contrast’s content by utilizing multiple reference contrasts in the reconstruction. Given number of contrasts , content could be defined as the local structure underlying all these contrasts, while style would encode intra-contrast variations of each individual contrast. This model would then require an (unpaired) image-domain training dataset of the contrasts. An open question is how to effectively scale the model and training complexity to contrasts. The PnP-CoSMo algorithm can subsequently be extended to utilize multiple reference contrasts to guide a given target contrast. The specific question to investigate then would be – given reference contrasts (where ), how to aggregate their contents such that the prior structural information about the target image is maximized?
@@
In terms of training data requirements, PnP-CoSMo can make use of the larger amounts of the available image-domain data and can be applied to situations where end-to-end methods are infeasible. In situations where end-to-end methods are indeed feasible, such as in our NYU fastMRI benchmark (Section 5.2), PnP-CoSMo was often superior to state-of-the-art methods in terms of both metrics and visual quality, especially at higher accelerations, e.g. at =5 with a difference of 0.01 SSIM, 0.04 HaarPSI, and 0.02 DISTS on average over MC-VarNet in Task-1 and of 0.01 SSIM, 0.08 HaarPSI, and 0.04 DISTS in Task-2. Additionally, PnP-CoSMo is, by design, immune to cross-contrast distribution shifts as opposed to end-to-end methods, which was demonstrated, e.g. in comparison with MC-VarNet which suffered a difference of up to 0.10 SSIM, 0.17 HaarPSI, and 0.06 DISTS on average when applied on out-of-distribution tasks. Moreover, unpaired image-domain pre-training of the content/style model boosts the practical applicability of PnP-CoSMo, e.g. in the case of LUMC-TRA where the T1W/T2W data imbalance was considerable. In our LUMC multi-coil benchmark, we limited our baselines to those algorithms that were feasible given the data constraint. As future work, a more comprehensive study should be conducted with a broader range of methods and evaluation criteria, e.g. comparing generalizability across accelerations and sampling patterns with multi-contrast unrolled networks, given a fixed budget of paired training data. Contrary to the training stage, we strictly assumed the spatial alignment of the reference and target images in the reconstruction stage. The CR module, which can implicitly correct minor registration errors, would likely break down at the typically observed levels of patient motion. A potential solution is to incorporate an online registration step to explicitly and efficiently correct for arbitrary inter-scan motion, thereby further improving practical applicability.

### v5

@@ 4.1.4 Preliminary radiological evaluation
Finally, as a supplement to the LUMC benchmark, we conducted a preliminary radiological evaluation with the goal of assessing the quality of PnP-CoSMo reconstructions in a concrete clinical context. The study was conducted on a small sample of the LUMC-TRA recon-test set. First, all 20 subjects from this set (see Table 2) were screened by a junior radiologist, for the presence of brain metastases, resulting in three scans included in this part of the study. The scoring was then performed independently by two senior neuroradiologists. The scoring system comprised two sets of criteria, namely visual quality and pathology. Four visual quality criteria were used, namely sharpness, noise, artifacts, and tissue contrast between gray and white matter and CSF. For pathology, we used three criteria, namely the number and sharpness of hyperintense areas within or surrounding metastases and the overall diagnostic quality of the scan for brain metastases. The images were scored using a five-point Likert scale (Mason et al., 2019) – (1) non-diagnostic, (2) poor, (3) fair, (4) good, and (5) excellent diagnostic quality.
Furthermore, we used PnP-CNN as a non-guided plug-and-play reconstruction baseline. We evaluated the two methods at four clinically realistic accelerations of , which are close to the routine clinical acceleration of =2 for this sequence. In total, nine images per patient, including the (24=8) reconstructions and the clinical ground truth (acquired at =2 and reconstructed by the vendor software), were presented to the two observers who scored each image individually, blinded to the reconstruction method and . The scores were then processed as follows. Given a reconstruction method and , the scores for each Likert-scale criterion were expressed as an aggregate by computing the mean across the three subjects and the two observers. The number of visible metastases was expressed as the fraction of total metastases across the three subjects. Moreover, the agreement between the two observers was measured using Spearman’s rank correlation coefficient, computed separately on the Likert-scale scores and the metastasis count.
@@ 4.2 Implementation details
We used a similar general residual architecture for content/style encoders and decoders as in the original MUNIT paper (Xuan et al., 2022), except with an additional layer at the end of the content encoder to produce content maps of the specified number of channels. We used 2 or 4 content channels, depending on the content downsampling factor. The discriminators were implemented as multi-scale PatchGAN networks enabling them to locally assess the input images for realism at different scales. We used 1 scale in the BrainWeb simulations and three scales in the NYU and LUMC benchmark. We additionally conditioned the discriminators on foreground masks of the images to penalize background signal, and we used spectral normalization (Miyato et al., 2018) to stabilize the GAN training. For hyperparameters, we found that pre-training loss weight values and paired fine-tuning values worked generally well for the in vivo NYU and LUMC datasets. Of all these hyperparameters, and , which control the disentanglement level, are most interesting. Therefore, the tuning of and for the synthetic BrainWeb data is presented later in Section 5.1.1, where the optimal value was found to be on this dataset. In the benchmarks, the step size of the CR module of PnP-CoSMo was tuned per acceleration for each dataset and the number of iterations was set to 200.
In the simulations, all content/style models were pre-trained for 200k and were not fine-tuned for the sake of simplicity, unless stated otherwise in Section 5.1. In both NYU and LUMC benchmarks, the content/style models were pre-trained for 400k iterations and fine-tuned for 50k iterations. We empirically found that a content/style model with content capacity =1 was optimal for the NYU dataset, whereas models with capacity =1/4 were optimal for the LUMC datasets. For the synthetic BrainWeb data, the effect of this parameter on the reconstructions is presented later in Section 5.1.2, where the optimal value was found to be =1 when both contrasts have equal spatial resolution.
All software was implemented in PyTorch, and image registration was performed using Elastix (Klein et al., 2009). All training runs were performed on a compute node with an NVIDIA Quadro RTX 6000 GPU.
@@ 5 Results
In the following sections we report the results of our extensive evaluation. In Section 5.1, we report our analyses on the synthetic BrainWeb dataset. These include a sensitivity analysis with respect to the level of disentanglement of content and style (Section 5.1.1), content capacity analysis exposing the relation between reference image resolution and the reconstruction quality (Section 5.1.2), convergence analysis of our iterative algorithm (Section 5.1.3), and a series of analyses on the CR component including evaluating its effectiveness in correcting for contrast-specific structure small misalignments and the influence of the CR step size (Section 5.1.4). In Section 5.2, we present the results of our benchmark against three end-to-end and three plug-and-play baselines on the public NYU fastMRI data. In Section 5.3, results of our benchmark on our in-house LUMC multi-coil raw data are reported. Given the real-world constraints of our clinical datasets, specifically a collection of partially-paired image-domain training data and complete lack of k-space training data, PnP-CoSMo is compared with classical unguided and guided reconstruction algorithms, plug-and-play unguided and guided methods, and image-to-image translation. Furthermore, our ablation study of PnP-CoSMo on these datasets demonstrates the contribution of the PFT and CR components of PnP-CoSMo. Finally, in Section 5.4, results of our small preliminary radiological evaluation are reported.
@@ 5.1.2 Content capacity analysis
PnP-CoSMo relies on a reference image of sufficiently high resolution to provide guidance to the reconstruction. The goal of this experiment was to determine the effect of lowering the reference image resolution on the reconstruction quality and using the content/style model’s content capacity (Eq. (14)) to explain this effect. We simulated 4 datasets, denoted as RefRes-, where represents the reference domain downsampling factor. In the =1 case, the T1W/T2W images had the same resolution (as that of the underlying tissue maps), whereas in the subsequent cases, T1W images were blurred to contain only the lower frequency components, while maintaining the same spatial resolution. For each dataset, we trained three content/style models with content capacity =, where is the content downsampling factor, which depends on the number of up/downsampling blocks in the networks. E.g. the =(1/2)2 model produced content maps half the spatial resolution of the =1 model.
@@
5.1.3 Convergence analysis
We explored the convergence of PnP-CoSMo by comparing it with two variants. The first used true content of the ground truth T2W image, thus assuming zero content discrepancy and representing an upperbound of PnP-CoSMo. The second used reference content with CR module disabled, representing a lowerbound of PnP-CoSMo where the non-zero content discrepancy is not corrected. Here and in the following simulation experiments, we used the RefRes-1 dataset and the =1 model, which was additionally fine-tuned on 2 training volumes for 50k iterations. Fig. 4 shows the convergence curves and intermediate reconstructions. The upperbound version, given , converged in a single iteration, while with , the ablated version converged equally fast but to a sub-optimal solution. Enabling the CR module closed the gap with the upperbound, although at the cost of slower convergence. It is also observed that the content discrepancy reduced when using CR and that the style estimate always converged within the first few iterations, as we hypothesized in Section 3.3.
5.1.4 Analyses on the content refinement procedure
The CR update step (Eq. 17) corrects of arbitrary discrepancies between the prior content and the measured k-space. We tested its robustness against two specific discrepancies: (a) contrast-specific structure (for e.g. a lesion) present in the target contrast but not in the reference contrast and (b) spatial misalignment in the reference contrast image. Additionally, the influence of the CR step size was also analyzed.
In order to test the effectiveness of the CR module in resolving structure present in the target contrast but absent in the reference contrast, we simulated a lesion in the T2W image and performed reconstruction with and without the CR module. As shown in Fig. 5, the CR module was able to recover the lesion fully at =2 and substantially at .
@@ 5.1.4 Analyses on the content refinement procedure
Finally, we also analyze the influence of the hyperparameter in Algorithm 1, which controls the CR step size. Fig. 7 shows reconstruction quality as a function of acceleration and k-space noise level, where we observe that the optimal decreased with the amount and quality of k-space data. This was expected as enforcing agreement with the k-space gradually becomes less beneficial, thereby contributing less to reconstruction quality.
@@ 5.2 Benchmark on public NYU fastMRI DICOM data
In the NYU benchmarking setup considered here, let the T2W reconstruction task be denoted as Task-1 and the T1W reconstruction task as Task-2. In Task-1, guided methods use the corresponding T1W image as reference, and in Task-2, the T2W images are used as reference. Unguided methods such as MoDL and PnP-CNN ignore this side-information. Fig. 8 shows evaluation metrics for PnP-CoSMo and the end-to-end baselines on these two guided reconstruction tasks.
We observe three main trends. First, comparing with the in-distribution reconstruction quality of the end-to-end baselines, PnP-CoSMo largely outperformed MoDL (0.05 throughout, except for SSIM at =2 and DISTS at on Task-1, and DISTS at =2 on Task-2) and MTrans (0.05 throughout), and was roughly comparable to MC-VarNet (e.g. with 0.05 on Task-2 in terms of DISTS at =3 and SSIM at =4) while often outperforming it (0.05 throughout, except SSIM at and DISTS at on Task-1, and SSIM at and DISTS at =2 on Task-2). Compared to the plug-and-play baselines, PnP-CoSMo largely outperformed PnP-CNN (0.05 throughout, except at =2) and was significantly superior to PnP-Diffusion and PROSIT (0.05 throughout for both). As shown in Fig. 9a, PnP-CoSMo could resolve certain fine details even at the highest acceleration of =5, which both end-to-end and plug-and-play baselines fell short at. In the pathological case shown in Fig. 9b, PnP-CoSMo produced sharper reconstructions while preserving the structure of the edema better than (the in-distribution) MC-VarNet, despite slightly lower metrics, when many features of the edema were absent in the reference image.
Second, considering out-of-distribution generalization, the end-to-end methods consistently dropped in performance on the out-of-distribution tasks (0.05 for all three models on both tasks at all 4 accelerations). This was expected since end-to-end models are explicitly trained to optimize for a given contrast task and will inevitably not generalize well across contrasts.
Third, PnP-CoSMo definitively outperformed the out-of-distribution variants of all the end-to-end models (0.05 throughout). Since PnP-CoSMo does not depend on any task-specific k-space training data owing to both its plug-and-play design and because of the bi-directionality of the content/style model it uses, it is not susceptible to the distribution shift of this kind. This can be observed in Fig. 10.
In summary, PnP-CoSMo produced generally equivalent or better reconstruction quality than the end-to-end methods despite its apparent handicap – i.e. completely ignoring the k-space training data and not being trained to explicitly solve the reconstruction problem. Moreover, this very feature of PnP-CoSMo makes it immune to contrast-related domain shifts. Finally, PnP-CoSMo was superior also to other plug-and-play methods, both unguided and guided ones, which share a similar design pattern and training data requirements.
@@ 5.4 Preliminary radiological evaluation
Fig. 13 plots the aggregated scores of our preliminary radiological evaluation and Fig. 14 shows a set of representative example reconstructions.
At , PnP-CoSMo produced reconstructions of at least ”fairly diagnostic” level (Likert score 3) of image sharpness, artifacts, tissue contrast, and overall diagnostic quality, while enabling the detection of at least 80% of all metastases. In terms of artifacts, PnP-CoSMo was consistently superior, as seen also in Fig. 14 (blue arrows). In terms of tissue contrast, it matched or exceeded PnP-CNN at and produced sharper reconstructions at (purple arrows in Fig. 14). However, in terms of noise, it was equal or slightly worse. The may be explained by the difference in nature between the denoiser-type model of PnP-CNN and the more complex content/style model of PnP-CoSMo, potentially leading to noise-like grainy texture in the latter which also interferes with image sharpness and tissue contrast. In terms of the pathology criteria, while both algorithms scored similarly on the metastasis count, PnP-CoSMo had slightly better diagnostic quality at =6. The general difference in reconstruction quality at =6 can be seen in Fig. 14 where, compared to PnP-CoSMo, the PnP-CNN reconstruction shows severe reconstruction artifacts, blurrier and more distorted features (purple arrows), and a poorly visible hyperintense area of a brain metastasis (salmon-pink arrows).
@@ 6 Discussion
Optimal content capacity (Section 3.2.3) was defined as a quantity representing the amount of shared local information contained in a two-contrast dataset. In terms of this quantity, we analyzed the effect of reference image resolution on reconstruction quality, obtaining a lowerbound for PnP-CoSMo to be effective (Section 5.1.2), specifically a reference resolution 1/8 of the target image resolution, corresponding to =1/16 of the total spatial information in the target. This quantity would depend on more fundamental factors such as MR sequence types, which we did not investigate here. While we limited our experiments to T1W and T2W sequences, PnP-CoSMo is, in principle, applicable to any pair of contrasts, with the image quality limited mainly by the amount of structural information shared between them. Future work could help empirically determine which sequence pairs are more amenable to guided reconstruction than others.
A consequence of our MUNIT-based model design was the model’s lack of dependence on the uni-directionality of the guided reconstruction problem, which allowed greater generalizability by design (Section 5.2). It is, in principle, also possible to model the content and style for more than two contrasts. This would enable not only the generalizability of our approach to any combinatorial set of guided reconstruction tasks, but also better estimation of the target contrast’s content by utilizing multiple reference contrasts in the reconstruction. Given number of contrasts , content could be defined as the local structure underlying all these contrasts, while style would encode intra-contrast variations of each individual contrast. This model would then require an (unpaired) image-domain training dataset of the contrasts. An open question is how to effectively scale the model and training complexity to contrasts. The PnP-CoSMo algorithm can subsequently be extended to utilize multiple reference contrasts to guide a given target contrast. The specific question to investigate then would be – given reference contrasts (where ), how to aggregate their contents such that the prior structural information about the target image is maximized?
@@
In terms of training data requirements, PnP-CoSMo can make use of the larger amounts of the available image-domain data and can be applied to situations where end-to-end methods are infeasible. In situations where end-to-end methods are indeed feasible, such as in our NYU fastMRI benchmark (Section 5.2), PnP-CoSMo was often superior to state-of-the-art unguided (MoDL) and guided (MC-VarNet and MTrans) networks in terms of both metrics and visual quality, especially at higher accelerations, e.g. at =5 with a difference of 0.01 SSIM, 0.04 HaarPSI, and 0.02 DISTS on average over MC-VarNet in Task-1 and of 0.01 SSIM, 0.08 HaarPSI, and 0.04 DISTS in Task-2, despite PnP-CoSMo completely ignoring the k-space training data and not being explicitly optimized for any reconstruction task in the training phase. Additionally, PnP-CoSMo is, by design, immune to cross-contrast distribution shifts as opposed to end-to-end methods, which was demonstrated, e.g. in comparison with MC-VarNet which suffered a difference of up to 0.10 SSIM, 0.17 HaarPSI, and 0.06 DISTS on average when applied on out-of-distribution tasks. Compared to other plug-and-play methods, which share a similar de-coupled design pattern and the image-only training data requirement as PnP-CoSMo, it produced sharper and higher-fidelity reconstructions and higher metrics, e.g. with a maximum difference of 0.05 SSIM, 0.13 HaarPSI, and 0.03 DISTS on average over PROSIT in Task-1 and of 0.03 SSIM, 0.10 HaarPSI, and 0.02 DISTS in Task-2. Moreover, the unpaired image-domain pre-training of the content/style model boosts the practical applicability of PnP-CoSMo, e.g. in the case of LUMC-TRA where the T1W/T2W image data imbalance was considerable (Table 2). In our LUMC multi-coil benchmark, we limited our baselines to those algorithms that were feasible given the data constraint. As future work, a more comprehensive study should be conducted with a broader range of methods and evaluation criteria, e.g. comparing generalizability across accelerations and sampling patterns with multi-contrast unrolled networks, given a fixed budget of paired training data. Contrary to the training stage, we strictly assumed the spatial alignment of the reference and target images in the reconstruction stage. The CR module, which can implicitly correct minor registration errors, would likely break down at the typically observed levels of patient motion. A potential solution is to incorporate an online registration step to explicitly and efficiently correct for arbitrary inter-scan motion, thereby further improving practical applicability.